# Media Coverage of War Victims: Journalistic Biases in Reporting on Israel and Gaza

### v1

@@ Results
With this data, we conducted our key analyses. We investigated how media outlets handled (de-)humanizing reporting, quantified the human costs of war, and reported on child victims. This allowed us to uncover nuanced media biases across different sources.
@@
After identifying doubt-casting phrases, we analyzed their occurrences for each side (Palestine vs. Israel) across four news sources. The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 915 CVN mentions, 98% of which target Palestinians—averaging over two CVNs doubted daily during the first year. This far exceeds CNN (145), NYT (131), and AJE (72). Common Source Doubting phrases like “Hamas-run health ministry” were heavily used by BBC, CNN, and NYT, whereas analogous expressions for Israel (such as “Knesset-run health ministry”) were absent. CNN and NYT also favored phrases like “Hamas-controlled enclave.” For Uncertainty in Numbers, terms such as “reportedly” and “claim” were predominantly used to describe Palestinian casualties by Western outlets, whereas AJE applied them more evenly across both sides (see Supplementary Tables 11–14 for detailed lists of phrases by outlet). Western sources attributed 71–90% of their cast-doubting phrases to doubting the sources of Palestinian CVNs, compared to 21% for AJE. Additionally, AJE was the only outlet to use Uncertainty in Numbers phrases in a balanced manner: 53% referred to Palestinians and 26% to Israelis, a proportion consistent with the fact that AJE published roughly twice as many articles about Palestinian CVNs as about Israeli ones. In contrast, the other outlets applied such language almost exclusively to Palestinians, with 3% or less referring to Israelis. These findings underscore in particular BBC’s significant reliance on doubt-casting phrases, particularly when reporting on Palestinian casualties, compared to other sources.
@@ Discussion
How did BBC, CCN, and NYT journalists manage to create the illusion of balance in the face of extreme disparities in military power and human suffering? The main strategy was to persistently refer back to the 7th of October massacre, even at a distance of months and in the absence of new events involving Israeli victims. Our analysis suggests that reiterations of October 7 were particularly pronounced during periods when significant events occurred on both sides, often overshadowing coverage of Palestinian suffering and diverting attention from the toll in Gaza. This strategy, which contradicts the common “agenda decay” phenomenon, not only creates an illusion of balance, but also reiterates the idea that the 7th of October was the starting point and “cause” of the subsequent events.

### v2

@@ Results
With this data, we conducted our key analyses. We investigated how media outlets handled (de-)humanizing reporting, quantified the human costs of war, and reported on child victims. In order to establish bias, we compared media reporting, wherever possible, with objective standards such as actual casualty rates, as explained in greater detail below. This allowed us to uncover nuanced media biases across different sources. Importantly, with the term ”bias”, we are not assuming any form of intentionality or malevolence, but describing observable asymmetries in coverage that reflect underlying reporting tendencies. The possible reasons for the occurrence of such biases will be discussed in later sections.
@@
It is important to note, however, that one would not necessarily expect a perfect match between actual casualty rates and casualty reporting, indeed, such correspondence was absent in all outlets, including AJE. Some degree of over-reporting of the events of October 7th compared to the following twelve months would be expected on the basis of well-documented psychological mechanisms. First, the October 7th massacre was extremely salient, involving a large number of victims in a short time span. Research on the availability heuristic shows that a single-mass casualty event tends to attract disproportionate attention and to be remembered more vividly than a higher, but gradually accumulating, death tolls [30]. Second, events that occur first in time or space tend to receive greater attention than those that follow (e.g., the ballot-order effect [31]. Thus, to the extent that the October 7 attack is perceived as the “starting point” of the conflict, it is understandable that it received over-proportional coverage. Nevertheless, such cognitive and temporal biases cannot account for the extent to which reporting equated casualties on both sides. For instance, while the 38 Israeli children killed on October 7th might be expected to draw more attention than the 38 Palestinian children killed, on average, each day over the subsequent year, it would be implausible for the cumulative 14,000 Palestinian child deaths to receive equivalent attention as those 38 Israeli deaths.
@@
After identifying doubt-casting phrases, we analyzed their occurrences for each side (Palestine vs. Israel) across four news sources. The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 915 CVN mentions, 98% of which target Palestinians, averaging over two CVNs doubted daily during the first year. This far exceeds CNN (145), NYT (131), and AJE (72). Common Source Doubting phrases like “Hamas-run health ministry” were heavily used by BBC, CNN, and NYT, whereas analogous expressions for Israel (such as “Knesset-run health ministry”) were absent. CNN and NYT also favored phrases like “Hamas-controlled enclave.” For Uncertainty in Numbers, terms such as “reportedly” and “claim” were predominantly used to describe Palestinian casualties by Western outlets, whereas AJE applied them more evenly across both sides (see Supplementary Tables 11–14 for detailed lists of phrases by outlet). Western sources attributed 71–90% of their cast-doubting phrases to doubting the sources of Palestinian CVNs, compared to 21% for AJE. Additionally, AJE was the only outlet to use Uncertainty in Numbers phrases in a balanced manner: 53% referred to Palestinians and 26% to Israelis, a proportion consistent with the fact that AJE published roughly twice as many articles about Palestinian CVNs as about Israeli ones. In contrast, the other outlets applied such language almost exclusively to Palestinians, with 3% or less referring to Israelis. These findings underscore in particular BBC’s significant reliance on doubt-casting phrases, particularly when reporting on Palestinian casualties, compared to other sources.
Doubting Palestinian sources is certainly justified, given the inherent difficulties in providing precise victim estimates during wartime. Statistics released by Palestinian authorities in Gaza were indeed frequently corrected and updated as new information became available. However, the same holds true for victim estimates on the Israeli side, where official statistics were revised repeatedly for up to six months after the events. Notably, skepticism toward Palestinian figures was often amplified by the practice of explicitly associating them with Hamas, an organization formally designated as terrorist group in the United States [77] and the United Kingdom [78], thereby highlighting their perceived lack of credibility. This rhetorical framing tends to obscure the fact that similar or identical casualty estimates were also reported and validated by independent researchers, international institutions, and humanitarian organizations (e.g., [79]).
@@ Discussion
How did BBC, CNN, and The NYT reporting convey an impression of (false) balance despite the extreme disparities in military power and human suffering? One contributing factor was the frequent return to the events of October 7th, even months later and in the absence of new incidents involving Israeli victims. Our analysis suggests that reiterations of October 7 were particularly pronounced during periods when significant events occurred on both sides, often overshadowing coverage of Palestinian suffering and diverting attention from the toll in Gaza. This practice, which contradicts the common “agenda decay” phenomenon, not only creates an illusion of balance, but also reiterates the idea that the 7th of October was the starting point and “cause” of the subsequent events. Importantly, we do not intend to suggest that the reiteration of October 7th was part of a deliberate strategy pursued by the analyzed media, however, even such unintentional biases may have contributed to the neglect of Palestians’ victimhood in public perception. Biases can arise from common and often automatic cognitive processes. As previously discussed, a single mass-casualty event tends to attract disproportionate attention and to be remembered more vividly than a higher but gradually accumulating death toll. In addition, events that occur first in time or space typically receive greater attention than those that follow. Together, these mechanisms may partly explain why the October 7th massacre remained a persistent reference point in subsequent reporting. At the same time, perfect parity in coverage should not be expected, given the profound asymmetry in casualties between the two sides and the unequal number of conflict-related events that could be reported.